Computability, Unsolvability, Randomness Math 497A: Student Research Projects

Stephen G. Simpson Due 8:00 AM, December 10, 2007

Dory Deines: unsolvability of the problem of integrability in finite terms. Bryan Gillespie: alternative notions of randomness. Jorge Gonzalez: Hilbert’s 10th Problem. Alison Green: randomness and convergence rates in the Strong Law of

Large Numbers. Sara Jensen: implementation of register machines in Mathematica. Chor Lam: effective Hausdorff dimension. Vincent Martinez: modular machines and unsolvability of the word prob-

lem for groups. Daniel McDonald: tilings of the plane. Timur Nezhmetdinov: finite lattices of Turing degrees. Matt Pancia: randomness and the Law of the Iterated Logarithm, etc. Noopur Pathak: randomness and the Lebesgue Differentiation Theorem. Peter Risch: alternative notions of randomness and low-for-randomness. Soumya Sanyal: mass problems. Stephanie Tougas: Kolmogorov complexity. Robin Tucker-Drob: creative sets and the Gödel Incompleteness Theo-

rems. Ashley Wheeler: finite lattices of Turing degrees and the jump operator.